OpenStudy (anonymous):

How to convert your exponential function P(x) in a logarithmic one and then into a base 10 logarithm.

OpenStudy (ranga):

Question is not clear. Can you clarify? "in a logarithmic one" is not clear.

OpenStudy (anonymous):

@ranga in logarithmic function

OpenStudy (ranga):

oh, "INTO a logarithmic function" not "in a logarithmic function" ?

OpenStudy (ranga):

\[\Large \log_{B} A = C \quad means \quad A = B^C\]

OpenStudy (ranga):

The first is in the logarithm form and the second is in the exponential form.

OpenStudy (anonymous):

haha yess sorry

OpenStudy (ranga):

\[\Large y = a^x\] is an exponential function. We can convert it to a log function by taking logarithm on both sides: \[\Large \log(y) = \log(a)^x = x * \log(a) = \log(a) * x\]

OpenStudy (ranga):

The log that we take can be logarithm to any base. But the ones most commonly used are logarithm to the base 10 and logarithm to the base e (called natural logarithm).